OpenStudy (anonymous):

The volume of a gas in a balloon is inversely proportional to the pressure on the gas from outside the balloon. If the volume of the gas is 6 cubic centimeters when the pressure on the gas is 18 kilograms per square centimeter, what is the volume when the pressure is 12 kilograms per square sentimeter?anyone can help me?

OpenStudy (anonymous):

Inversely proportional means k over (divided by) so V = k/pressure.. Substituting you get 6 = k/18 if you write this 6/1 = k/18 When you have a fraction = fraction you cross multiply. So you get 6*18 = 1*k so 108 = k.. Now you go to the second part V = k/pressure. V = we don't know. k = 108... pressure = 12 so V = 108/12 so Volume = 9 cubic centimeters